# Efficient Epistemic Uncertainty Estimation in Regression Ensemble Models Using Pairwise-Distance Estimators

### v3

@@ 4.1 Nflows Base
NFs have been classically applied to unsupervised tasks (Tabak & Vanden-Eijnden, 2010; Tabak & Turner, 2013; Rezende & Mohamed, 2015), though NFs have been adapted to a supervised learning setting for regression tasks (Winkler et al., 2019; Ardizzone et al., 2019). Using the structure of Winkler et al. (2019), one can define a supervised NF as,
$p_{y|x}(y|x)$ $=p_{b|x,\theta}(g^{-1}_{\phi}(y,x))$
×$\times|\det(J(g^{-1}_{\phi}(y,x)))|,$
$\log(p_{y|x}(y|x))$ $=\log(p_{b|x,\theta}(g^{-1}_{\phi}(y,x)))$
$+\log(|\det(J(g^{-1}_{\phi}(y,x)))|),$
where $p_{y|x}$ is the output distribution, $p_{b|x,\theta}$ is the base distribution with parameters $\theta$, $J$ refers to the Jacobian, and $g^{-1}_{\phi}:y\times x\mapsto b$ is the bijective mapping with parameters $\phi$. For a more comprehensive review of NFs, refer to Papamakarios et al. (2021). Nflows Base creates an ensemble in the base distribution,
@@
where $p_{b|x,\theta}(b|x,\theta_{j})=N(\mu_{\theta_{j}}(x),\Sigma_{\theta_{j}}(x))$, $\mu_{\theta_{j}}(x)$ and $\Sigma_{\theta_{j}}(x)$ denote the mean and covariance for input $x$ and conditioned on $\theta_{j}$. These parameters are modeled using a neural network with fixed dropout masks to create an ensemble, and diversity is created by randomization and bootstrapping. By constructing the ensemble within the base distribution, one can make use of closed-form pairwise-distance formulae as Berry & Meger (2023) showed that estimating epistemic uncertainty in the base distribution is equivalent to estimating it in the output distribution.
@@
Exploiting these closed-form pairwise-distance formulae, Berry & Meger (2023) showed that Nflows Base outperforms naive NF ensemble methods when estimating epistemic uncertainty. The aleatoric uncertainty from Equation 1 can be estimated in the base distribution space. Therefore aleatoric uncertainty can be computed analytically. This does not apply to the other quantity of Equation 1, total uncertainty, and thus MC techniques were applied in the past to estimate epistemic uncertainty.
5 Epistemic Uncertainty Estimation with PaiDEs
Our extension to prior methodologies empowers the estimation of epistemic uncertainty, the mutual information between $y$ and $\theta$, without the need for sampling. We illustrate this capability for Normalizing Flows (NFs) in the main paper and for probabilistic network ensembles in Figure 10 (Lakshminarayanan et al., 2017). Note that our proposed estimators can be applied to any ensemble model whose component output distributions have closed-form pairwise-distances. However, given that the space of models with these properties is large, we have decided to focus on Nflows Base for its expressive properties and PNEs as they are widely used (Lakshminarayanan et al., 2017; Chua et al., 2018; Berry & Meger, 2023).
5.1 Epistemic Uncertainty Estimators
$\hat{I}_{\rho}(y,\theta)$ $=\hat{H}_{\rho}(y|x)-E_{p(\theta)}\left[H(y|x,\theta)\right]$
$=-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-D(p_{i}\;\|\;p_{j})\right)},$ (11)
as $E_{p(\theta)}\left[H(y|x,\theta)\right]=H(y|x,\theta)$ leading to cancellation. PaiDEs provide a succinct estimator that can capture epistemic uncertainty with only the pairwise distances between components, thus eliminating reliance on sample-based techniques. We propose the following specific estimators:
$\hat{I}_{Bhatt}(y,\theta)$ $=-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-D_{Bhatt}(p_{i}\;\|\;p_{j})\right)},$
where $D_{Bhatt}(p_{i}\;\|\;p_{j})$ and $D_{KL}(p_{i}\;\|\;p_{j})$ are defined for Gaussians in Appendix B.
5.2 Combination of PaiDEs & Nflows Base
@@
By combining Nflows Base and PaiDEs, we construct an expressive non-parametric model capable of capturing intricate aleatoric uncertainty in the output distribution while efficiently estimating epistemic uncertainty in the base distribution. Unlike previously proposed methods, we are able to estimate epistemic uncertainty without taking a single sample. Figure 6 in the Appendix shows an example of the distributional pairs that need to be considered in order to estimate epistemic uncertainty for an Nflows Base model.
@@ 6 Experimental Results
[…]
Figure 2: In the right graphs, the brown dots are sampled from Nflows Base and the cyan dots are the ground-truth data. The left graphs depicts the epistemic uncertainty estimates corresponding to our two proposed estimators and the density of the ground-truth data in the green histogram.
To evaluate our method, we tested each PaiDE (KL $\hat{I}_{KL}(y,\theta)$ and Bhatt $\hat{I}_{Bhatt}(y,\theta)$) on two 1D environments, as has been previously proposed in the literature (Depeweg et al., 2018). Additionally, we present 4 multi-dimensional environments. In contrast to previous papers (Postels et al., 2020; Ash et al., 2021), we evaluate each method on high dimensional output space (up to 257) to demonstrate the utility of PaiDEs in this setting. Note that, for all experiments, the model components are assumed to be uniform, $\pi_{j}=\frac{1}{M}$, independent of $x$. All model hyper-parameters are contained in Appendix B and the code can be found in the supplementary material.
@@ 6.1 Data
BatchBALD 6.62 $\pm$ 1.84 6.07 $\pm$ 0.12 0.06 $\pm$ 0.03 0.62 $\pm$ 0.18 1.93 $\pm$ 0.3 11.28 $\pm$ 0.16
BADGE 6.03 $\pm$ 1.68 6.25 $\pm$ 0.33 0.07 $\pm$ 0.04 0.62 $\pm$ 0.21 2.11 $\pm$ 0.26 10.99 $\pm$ 0.33
BAIT 7.39 $\pm$ 0.58 6.2 $\pm$ 0.41 0.08 $\pm$ 0.05 0.68 $\pm$ 0.23 1.91 $\pm$ 0.22 11.01 $\pm$ 0.23
LCMD 5.69 $\pm$ 1.03 6.24 $\pm$ 0.34 0.06 $\pm$ 0.03 0.64 $\pm$ 0.2 1.84 $\pm$ 0.17 11.35 $\pm$ 0.16
MC 1.54 $\pm$ 0.17 6.01 $\pm$ 0.04 0.06 $\pm$ 0.01 0.32 $\pm$ 0.02 1.01$\pm$ 0.05 10.71 $\pm$ 0.43
KL 1.47$\pm$ 0.15 6.01 $\pm$ 0.04 0.05$\pm$ 0.04 0.3 $\pm$ 0.03 1.06 $\pm$ 0.08 3.4 $\pm$ 0.48
Bhatt 1.55 $\pm$ 0.31 6.0$\pm$ 0.04 0.05$\pm$ 0.01 0.29$\pm$ 0.03 1.1 $\pm$ 0.07 3.37$\pm$ 0.44
@@
We evaluated PaiDEs on two 1D benchmarks, hetero and bimodal. The ground-truth data for hetero and bimodal can be seen in Figure 2 on the right graphs with the cyan dots. For hetero, there are two regions with low density (2 and -2) as can be seen by the green bar chart in the left graph which corresponds to the density of the ground-truth data. In these regions, one would expect a model to have high epistemic uncertainty. For bimodal, the number of data points drops off as x increases, thus we would expect a model to have epistemic uncertainty grow as x does. All details for data generation are contained in Appendix C.
In addition to the 1D environments, we tested our methods on four multi-dimensional environments: Pendulum, Hopper, Ant, and Humanoid (Todorov et al., 2012). Replay buffers were collected and the dynamics for each environment was modeled, $f_{\theta}(s_{t},a_{t})=\hat{s}_{t+1}$. The choice of multi-dimensional environments is motivated by their common use as benchmarks and their higher-dimensional output space, providing a robust validation of our methods. Note that, for Ant and Humanoid, the dimensions representing their contact forces were eliminated due to a bug in Mujoco-v2111More information can be found here: https://github.com/openai/gym/issues/1541..
6.2 1D Experiments
Figure 3: RMSE on the test set at the $100^{th}$ acquisition batch of the MC estimator on the Hopper environment as the number of samples, $K$, per input, $x$, increases. Experiment were run across 10 seeds and the mean is being reported.
Our 1D environments provide empirical proof that PaiDEs can accurately measure epistemic uncertainty. Figure 2 depicts that both KL and Bhatt are proficient at estimating the epistemic uncertainty as each method shows an increase in epistemic uncertainty around 2 and -2 on the hetero setting. This can be seen from the blue and orange lines with both estimators performing similarly.
A similar pattern can be seen for the bimodal setting in Figure 2, which shows that both Bhatt and KL can accurately capture epistemic uncertainty. Each estimator shows the pattern of increasing epistemic uncertainty where the data is more scarce. Both examples show accurate epistemic uncertainty estimation with no loss in aleatoric uncertainty representation, as demonstrated in the right graphs in Figure 2: the brown dots closely match the cyan dots. Note that for visual clarity, the epistemic uncertainty has been scaled to 0-1, while ensuring that the relative properties of the estimators are preserved. Further discussion on over- and underestimation is provided in Appendix D.
@@ 6.3 Active Learning
Figure 4: On the left, the amount of time taken for each estimator across the different settings (1, 3, 11, 27, 257 dimensions). On the right, the amount of time taken for PaiDEs as the number of ensemble components increases for the 257 dimensional setting. Results are averaged over 10 seeds and shown on a log scale.
@@
While the 1D experiments provide evidence of PaiDEs’ effectiveness for estimating epistemic uncertainty, the active learning experiments extend this evaluation to higher-dimensional data and for a decision-making task. In order to benchmark our method, we compare against four state-of-the-art active learning frameworks: BatchBALD (Kirsch et al., 2019), BADGE (Ash et al., 2019), BAIT (Ash et al., 2021) and LCMD (Holzmüller et al., 2023). Additionally, a random baseline is included. The training set is initialized with 100 or 200 data points for 1D and multi-dimensional environments, respectively. In each acquisition batch, 10 points are added. For the Monte Carlo (MC) estimator, 1,000 unseen inputs were sampled, and their epistemic uncertainties estimated, except for the Humanoid environment where only 100 new inputs were sampled due to computational constraints. In contrast, PaiDEs sampled 10,000 new inputs and estimated their epistemic uncertainties for each environment including Humanoid. This underscores one advantage of PaiDEs over MC estimators, as they can estimate epistemic uncertainty over larger regions at a lower computational cost than their MC counterparts. Note that the other benchmarks each sampled 10,000 inputs as well and their respective acquisition functions applied. The root mean squared error (RMSE) on the test set was calculated at each acquisition batch to evaluate performance.
Table 1 displays the performance of each framework on the $100^{th}$ acquisition batch. In each environment, we conducted a Welch’s t-test that compares both PaiDE estimators against each baseline. We included a Holm–Bonferroni correction to control the family-wise error rate, for more information refer to Appendix G. For each data setting, the PaiDEs reach lower or comparable RMSEs compared to each benchmark, demonstrating their efficacy in estimating epistemic uncertainty. Notably, in higher dimensions, PaiDEs outperform other methods in a statistically significant manner, as are shown in the Humanoid. The learning curves can be seen in Figure 5.
@@
To conduct a more in-depth analysis of our proposed method, we compared PaiDEs to MC estimators with a varying sample size, $K$, in the Hopper environment. We expected that MC estimators would perform on par with PaiDEs with a sufficient number of samples. This trend is illustrated in Figure 3. The MC estimator does reach the peak performance of PaiDEs given enough samples but does not perform better than PaiDEs, suggesting that PaiDEs are sufficient when estimating epistemic uncertainty in this context.
6.4 Time Analysis & Limitations
Figure 5: Mean RMSE on the test set as data was added to the training set for Nflows Base.
@@ 6.4 Time Analysis & Limitations
In addition to benchmarking PaiDEs on active learning experiments, we provide an analysis of the time gains across our experiments. The left hand side of Figure 4 depicts the speed increase that can be gained using PaiDEs over an MC approach. A 1-2 order magnitude of improvement can be seen. The estimates are obtained from the active learning experiments, and the number of dimensions correspond to each of the environments.
A weakness of PaiDEs is that, as the number of components increases, the computational cost rises due to its quadratic complexity. In the instance where the distance is not symmetric, KL-divergence, $M^{2}-M$ pairwise-distances need to be computed. For symmetric distances, such as Bhattacharyaa distance, only $\frac{M^{2}-M}{2}$ distances need to be computed. The right hand side of Figure 4 shows an analysis of the time taken as the number of ensembles grow. Note that for Bhatt, the time costs could be improved upon using the symmetry logic described, the results shown calculated all pairwise-distances. Despite the growing complexity of PaiDEs with the number of components, this is normally not a problem for deep learning ensembles as they have a relatively low number (5-10) of components (Osband et al., 2016; Chua et al., 2018).
An additional limitation is the bias introduced by PaiDEs, which MC estimators do not suffer from. It is essential to note that, in the context of active learning, epistemic uncertainty serves as a relative quantity for comparing potential acquisition points. The introduction of bias from PaiDEs does not impact the relative relationship of epistemic uncertainty between different data points in our experiments. We demonstrate that the relative relationship of epistemic uncertainty remains intact in Appendix D.

### v4

@@ 4.2 Nflows Base
In this study, we utilize an ensemble technique known as Nflows Base, which has previously demonstrated robust performance in estimating both aleatoric and epistemic uncertainty on simulated robotic datasets by leveraging NFs to create ensembles [4].
NFs have traditionally been applied to unsupervised tasks [54, 53]. However, NFs have also been adapted for supervised learning tasks, particularly for regression [59, 1]. Using the structure described in Winkler et al. [59], a supervised NF is defined as:
$p_{y|x}(y|x)$ $=p_{b|x,\theta}(g^{-1}_{\phi}(y,x))|\det(J(g^{-1}_{\phi}(y,x)))|,$
$\log(p_{y|x}(y|x))$ $=\log(p_{b|x,\theta}(g^{-1}_{\phi}(y,x)))+\log(|\det(J(g^{-1}_{\phi}(y,x)))|),$
where $p_{y|x}$ is the output distribution, $p_{b|x,\theta}$ is the base distribution with parameters $\theta$, $J$ refers to the Jacobian, and $g^{-1}_{\phi}:y\times x\mapsto b$ is the bijective mapping with parameters $\phi$. For a complete review of NFs, refer to Papamakarios et al. [47]. Nflows Base creates an ensemble in the base distribution,
@@
where $p_{b|x,\theta_{j}}(b|x,\theta_{j})=N(\mu_{\theta_{j}}(x),\Sigma_{\theta_{j}}(x))$, $\mu_{\theta_{j}}(x)$ and $\Sigma_{\theta_{j}}(x)$ denote the mean and covariance for input $x$ and conditioned on $\theta_{j}$. To encourage diversity in the ensemble, each member is initialized with different random weights, trained on a bootstrapped subset of the data, and assigned a fixed dropout mask sampled at the start of training (with $p=0.5$), similar to [1]. This mask remains constant throughout training and inference. By constructing the ensemble within the base distribution, one can make use of closed-form pairwise-distance formulae as Berry and Meger [4] showed that estimating epistemic uncertainty in the base distribution is equivalent to estimating it in the output distribution.
@@
Exploiting these closed-form pairwise distance formulae, Berry and Meger [4] showed that Nflows Base outperforms naive NF ensemble methods when estimating epistemic uncertainty. The aleatoric uncertainty from Equation 1 can be estimated in the base distribution space, and therefore can be computed analytically. However, this approach does not extend to the total uncertainty in Equation 1, which is why MC techniques have traditionally been used to estimate epistemic uncertainty.
4.3 Integrating Pairwise Epistemic Estimators with Nflows Base Ensembles
@@
By combining Nflows Base and PairEpEsts, we construct an expressive non-parametric model capable of capturing intricate aleatoric uncertainty in the output distribution while efficiently estimating epistemic uncertainty in the base distribution. Equation 9 then becomes,
$\hat{I}_{\rho}(y,\theta)=-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-D_{\rho}(p_{b|x,\theta_{j}}\;\|\;p_{b|x,\theta_{i}})\right)}.$ (10)
Unlike previously proposed methods, we are able to estimate epistemic uncertainty without taking a single sample. Figure 7 in the Appendix shows an example of the distributional pairs that need to be considered in order to estimate epistemic uncertainty for an Nflows Base model. By applying PairEpEsts in the base distribution space we are able to capture epistemic uncertainty in the output space [4]. This approach enables the use of established formulae for computing distributional distances.
PairEpEsts can in principle be extended to skewed or heavy-tailed base distributions, provided that the pairwise premetric admits a closed-form expression. Recent work on tail-adaptive normalizing flows [40, 25] offers a promising direction for incorporating heavy-tailed behavior into normalizing flows. Integrating these approaches with PairEpEsts represents an important avenue for future research, particularly in domains where tail risk is critical.
4.4 Integrating Pairwise Epistemic Estimators with Probabilistic Network Ensembles
In addition to Nflows Base ensembles, we extend PairEpEsts to PNEs, which are commonly used for uncertainty quantification in regression [12, 38, 39]. PNEs comprise multiple components that produce Gaussian predictive distributions with learned means and variances.
Applying PairEpEsts to PNEs leverages the closed-form pairwise distances between these Gaussian outputs to estimate epistemic uncertainty without sampling. Formally, the estimator is:
$\hat{I}_{\rho}(y,\theta)=-\sum_{i=1}^{M}\pi_{i}\ln\sum_{j=1}^{M}\pi_{j}\exp\bigl(-D_{\rho}(p_{y|x,\theta_{j}}\parallel p_{y|x,\theta_{i}})\bigr),$
where $p_{y|x,\theta_{i}}$ denotes the $i$-th output distribution. Our experiments in Appendix E show that PairEpEsts with PNEs achieve comparable or better epistemic uncertainty estimates than baselines, especially in high-dimensional settings. Although PNEs are less expressive than normalizing flow ensembles, their combination with PairEpEsts offers a practical and efficient framework for scalable uncertainty quantification across diverse ensemble models.
@@ 5 Experimental Results
(b) bimodal
Figure 2: In the right graphs, the brown dots are sampled from Nflows Base and the cyan dots are the ground-truth data. The left graphs depict the epistemic uncertainty estimates corresponding to our two proposed estimators and the density of the ground-truth data in the green histogram.
To evaluate our method, we tested each of our PairEpEsts (KL and Bhatt) on two 1D environments, as has been previously proposed in the literature [15]. Additionally, we present 4 multi-dimensional environments. In contrast to previous papers [48, 2], we evaluate each method on high dimensional output space (up to 257) to demonstrate the utility of PairEpEsts in this setting. Note that, for all experiments, the model components are assumed to be uniform, $\pi_{j}=\frac{1}{M}$, independent of $x$. All model hyper-parameters are contained in Appendix B and the code can be found in the supplementary material. The experimental design follow previous literature [4].
@@ 5.1 Data
We evaluated our estimators on two 1D benchmarks, hetero and bimodal. The ground-truth data for hetero and bimodal can be seen in Figure 2 on the right graphs with the cyan dots. For hetero, there are two regions with low density (2 and -2) as can be seen by the green bar chart in the left graph which corresponds to the density of the ground-truth data. In these regions, one would expect a model to have high epistemic uncertainty. For bimodal, the number of data points drops off as x increases, thus we would expect a model to have epistemic uncertainty grow as x does. All details for data generation are contained in Appendix C.
In addition to the 1D environments, we tested our methods on four multi-dimensional environments: Pendulum, Hopper, Ant, and Humanoid [55]. Replay buffers were collected and the dynamics for each environment was modeled, $f_{\theta}(s_{t},a_{t})=\hat{s}_{t+1}$. The choice of multi-dimensional environments is motivated by their common use as benchmarks and their higher-dimensional output space, providing a robust validation of our methods. Note that, for Ant and Humanoid, the dimensions representing their contact forces were eliminated due to a bug in Mujoco-v2.
In the Mujoco environment, the input consists of the current state and action, which are used to predict the next state based on the dynamics of the environment. In the hetero- and bimodal environments, we simplify the task to take a single input variable x, and predict the output variable y.
5.2 1D Experiments
@@
BatchBALD 1.54 $\pm$ 0.16 6.42 $\pm$ 0.65 0.13 $\pm$ 0.05 0.87 $\pm$ 0.27 0.94 $\pm$ 0.03 5.33 $\pm$ 1.2
[…]
BAIT 1.51 $\pm$ 0.14 6.26 $\pm$ 0.33 0.17 $\pm$ 0.05 1.06 $\pm$ 0.5 0.94 $\pm$ 0.04 11.01 $\pm$ 0.23
MC (BALD) 1.54 $\pm$ 0.17 6.01 $\pm$ 0.04 0.06 $\pm$ 0.01 0.32 $\pm$ 0.02 1.01 $\pm$ 0.05 10.71 $\pm$ 0.43
[…]
@@
Our 1D environments provide empirical proof that PairEpEsts can accurately measure epistemic uncertainty. Figure 2 depicts that both PairEpEsts are proficient at estimating the epistemic uncertainty as each method shows an increase in epistemic uncertainty around 2 and -2 on the hetero setting. This can be seen from the blue and orange lines with both estimators performing similarly.
A similar pattern can be seen for the bimodal setting in Figure 2, which shows that both PairEpEsts can accurately capture epistemic uncertainty. Each estimator shows the pattern of increasing epistemic uncertainty where the data is more scarce. Both examples show accurate epistemic uncertainty estimation with no loss in aleatoric uncertainty representation, as demonstrated in the right graphs in Figure 2: the brown dots closely match the cyan dots. Note that for visual clarity, the epistemic uncertainty has been scaled to 0-1, while ensuring that the relative properties of the estimators are preserved. Further discussion on over- and underestimation is provided in Section 6. While simple count-based heuristics may suffice in toy 1D settings like Figure 2, they fail to scale to higher-dimensional spaces where data is sparse and frequency statistics are uninformative. Moreover, they cannot exploit learned representations of the model, which are essential for structured, high-dimensional control tasks such as Hopper, Ant, and Humanoid. For these reasons, principled estimators such as PairEpEsts are required in realistic settings.
@@ 5.3 Active Learning
While the 1D experiments provide evidence of our estimators’ effectiveness for estimating epistemic uncertainty, the active learning experiments extend this evaluation to higher-dimensional data and for a decision-making task. In order to benchmark our method, we compare against four state-of-the-art active learning frameworks: BatchBALD which maximizes information gain over batches using MC approximations [34], BADGE which uses gradient-based selection to identify the most informative data points for model improvement [3] and BAIT selects data based on uncertainty estimation through Bayesian active learning [2]. Additionally, a random baseline is included. The training set is initialized with 100 or 200 data points for 1D and multi-dimensional environments, respectively. In each acquisition batch, 10 points are added. For the MC estimator, $10^{3}$ candidate inputs were sampled to construct each acquisition batch, with their epistemic uncertainties estimated using $K=5000$. In the Humanoid environment, only 100 candidates were used due to computational constraints. In contrast, PairEpEsts sampled $10^{4}$ candidate inputs and estimated their epistemic uncertainties for each environment including Humanoid. This underscores one advantage of our estimators over MC estimators, as they can estimate epistemic uncertainty over larger regions at a lower computational cost than their MC counterparts. Note that the other benchmarks each sampled $10^{4}$ inputs as well and their respective acquisition functions applied. The RMSE on the test set was calculated at each acquisition batch.
Figure 3: Mean RMSE on the test set as additional data is incorporated into the training set for Nflows Base. Both proposed methods perform comparably or significantly better than baselines, particularly in high-dimensional settings.
Table 1 shows the performance of each framework on the 100th acquisition batch. Welch’s t-test, with a Holm–Bonferroni correction (see Appendix G), compares our estimators to baselines in each environment. PairEpEsts achieve lower or comparable RMSEs, demonstrating their efficacy in estimating epistemic uncertainty. In higher dimensions, our estimators significantly outperform other methods, as shown in Humanoid. Learning curves are presented in Figure 3. Note that in the bimodal setting, all methods converge to similar performance and are indistinguishable in the plot. BatchBALD and BADGE, originally designed for classification, required adaptation for regression, facing challenges due to differences in uncertainty measure computation. Similarly, BAIT, designed for 1D regression, did not perform well in high-dimensional settings. The dashed line indicates the performance of our model when trained on the full dataset, serving as an upper bound for comparison.
Figure 4: RMSE on the test set at the $100^{th}$ acquisition batch of the MC estimator on the Hopper environment for Nflows Base as the number of samples, $K$, per input, $x$, increases. Experiment were run across 10 seeds and the mean and stdev is being reported.
@@
For a deeper analysis of our method, we compared our estimators to MC estimators with a varying sample size, $K$, in the Hopper environment. As expected, MC estimators perform on par with PairEpEsts with a sufficient number of samples. This trend is illustrated in Figure 4. The MC estimator reaches the peak performance of our estimators given enough samples but does not perform better than PairEpEsts, suggesting that our estimators are sufficient to estimate epistemic uncertainty.
5.4 Time Analysis
@@
In addition to benchmarking our estimators on active learning experiments, we provide an analysis of the time gains across our experiments. Figure 6 depicts the speed increase that can be gained using PairEpEsts over an MC approach. PairEpEsts provide a 1–2 order of magnitude runtime improvement over Monte Carlo approaches. The estimates are obtained from the active learning experiments.
6 Limitations
PairEpEsts have quadratic computational cost as the number of components grows. For asymmetric distances like KL-divergence, $M^{2}-M$ pairwise distances are computed, while symmetric distances like Bhattacharyya require only $\frac{M^{2}-M}{2}$. Figure 6 shows timing as ensemble size increases. Bhattacharyya’s cost could be further reduced using symmetry. Despite this, deep learning ensembles usually have few components (5–10) [45, 12], making the complexity manageable. The limitation is more relevant in large ensembles, such as weather forecasting [57].
PairEpEsts introduce a small bias absent in MC estimators. However, since active learning depends on the relative ordering of acquisition scores, this bias does not affect performance. As shown in Appendix D, the rankings are well preserved (Table 5). The relationships are illustrated in Figure 8, with epistemic uncertainty values normalized between 0 and 1 in Figure 2.